# Pinching-Antenna Systems (PASS): Architecture Designs, Opportunities, and Outlook

### v2

@@ Abstract
This article proposes a novel design for the Pinching Antenna SyStems (PASS) and advocates simple yet efficient wireless communications over the “last meter”. First, the potential benefits of PASS are discussed in the paper by reviewing an existing prototype. Then, the fundamentals of PASS are introduced, including their physical principles, signal models, and communication designs. In contrast to existing multi-antenna systems, PASS brings a novel concept, termed Pinching Beamforming, which is achieved by dynamically adjusting the positions of pinching antennas (PAs). Based on this concept, a couple of practical transmission architectures are proposed for realizing the full potentia of PASS, namely non-multiplexing and multiplexing architectures. More particularly, 1) the non-multiplexing architecture is featured by simple baseband signal processing and relies on the pinching beamforming only; while 2) the multiplexing architecture provides enhanced signal processing capabilities with joint baseband and pinching beamforming, which can be further divided into sub-connected, fully-connected, and phase-shifter-based fully-connected schemes. Furthermore, several emerging scenarios are put forward for integrating PASS into future wireless networks. As a further advance, by demonstrating a few numerical case studies, the significant performance gain of PASS is revealed compared to conventional multi-antenna systems. Finally, several research opportunities and open problems of PASS are highlighted.
@@ I Introduction
Building on this momentum, recent advancements in intelligent and flexible antenna designs have further introduced novel approaches for channel reconfiguration. Technologies, such as reconfigurable intelligent surfaces [3, 4], dynamic metasurface antennas [5], fluid antennas [6], and movable antennas [7], can dynamically adjust their positions, apertures, and electromagnetic properties, enabling precise control of wireless channel conditions. By optimizing their reconfigurable parameters, these novel antenna types and array architectures offer flexibility in controlling the effective end-to-end channel gain. This capability has demonstrated great potential for improving system performance and has attracted considerable research interests.
Despite their potentials to enhance network throughput, existing flexible-antenna systems face certain limitations. Firstly, those existing antenna reconfigurations are often confined to apertures spanning only a few wavelengths, which limits their ability to mitigate large-scale path loss. Secondly, in existing reconfigurable systems, it is costly, if not impossible, to change their antenna configuration, i.e., add or remove antennas once these systems are established. To address these challenges, a more flexible yet efficient new design—Pinching-Antenna SyStems (PASS) come into researchers’ sight [8, 9].
@@
I-A The Road to Near-Wired: A Prototype of PASS
The first prototype of PASS was developed and demonstrated by NTT DOCOMO in 2021 [8], as illustrated in Figure 1. A PASS consists of a specialized transmission line known as a dielectric waveguide. Similar to leaky-wave antennas, electromagnetic waves within the waveguide are emitted by pinching small separate dielectric particles at specific points along its length. These dielectrics are typically attached to the tips of plastic clips or clothespins, forming structures known as pinching antennas (PAs). By leveraging the waveguide’s ability to transmit or receive radio waves at any pinched point, PASS enables wireless communication precisely where needed [8].
The waveguide can be pinched at multiple points to provide coverage in specific locations, allowing for flexible and dynamic antenna deployment. Signal emission at a particular location ceases immediately upon removing the pinching antenna, making the process as simple as adding or removing a clothespin from a clothesline, as illustrated in Figure 1. This adaptability enables PASS to establish communication zones tailored to users’ locations and environmental conditions, offering exceptional flexibility in antenna deployment [8].
Unlike conventional flexible-antenna systems, the length of a PASS waveguide can be extended arbitrarily, allowing PAs to be positioned closer to users to establish strong line-of-sight (LoS) links [9]. This adaptability effectively creates a near-wired connection, ensuring favorable channel conditions and shifting the focus of wireless communication networks from the “last mile” to the more manageable “last meter”. This unique capability is reflected in the name “PASS”, highlighting its ability to provide the shortest and most efficient wireless connections between the PAs and the wireless transceivers, which effectively pass wireless signals with minimal loss.
@@ I-B Key Advantages for Employing PASS in Wireless Communications
  • Enhanced LoS Stability and “Last Meter” Communications: By extending the waveguide and enabling the flexible deployment of PAs, PASS effectively mitigate large-scale path loss and establish stable LoS communication environments. This capability enhances “last meter” communications and expands network coverage, which is not achievable with other flexible-antenna technologies.
@@
  • Cost-Effective and Scalable Implementation: PASS are simple to implement and economical to deploy, not only because of the low cost of waveguides, but also because PASS only require adding or removing dielectric materials at selected points along the waveguide [8], which makes PASS a scalable choice for practical network design.
These advantages have fueled growing research interest in PASS. The performance of PASS has been analyzed in terms of its ergodic rate [9] and its array gain [11]. Furthermore, various activation algorithms have been proposed to optimize the locations of PAs along a single waveguide; see [9] and relevant references. However, despite these initial efforts, the design of PASS for wireless communications remains in its early stages. This gap motivates us to provide a systematic introduction to PASS. In this article, we first discuss how PASS differ from conventional multi-antenna systems. We then propose several novel beamforming architectures to facilitate the design of pinching beamforming. This is followed by an exploration of promising application scenarios for PASS and an evaluation of their performance. Finally, we highlight the open challenges of PASS to inspire future research in this promising research area.
@@ II Pinching Antenna Fundamentals:
In this section, we introduce the fundamentals of PAs and highlight the basic differences from the conventional electronic antennas from their physical principle, signal model, and communication system design perspective, respectively.
@@ II-A Physical Principle
Unlike conventional electronic antennas which rely on electronic devices to convert alternating electric currents into electromagnetic (EM) waves, the physical principle of PAs is based on the phenomenon of EM coupling that arises when a separate dielectric is brought into close proximity with a dielectric waveguide [8]. This coupling enables the power exchange between the waveguide and the separate dielectric, as illustrated in the bottom-right of Fig. 1. Therefore, by pinching a separate dielectric onto the waveguide, the EM waves originally confined within the waveguide are coupled into the separate dielectric, where they are subsequently radiated into free space, thereby forming an “antenna”. Unlike conventional electronic antennas, PAs are entirely passive, simpler to manufacture and deploy, and hence provide greater position flexibility.
@@
$P_{\mathrm{guide}}$ $=1-F\sin^{2}(\kappa L),\quad P_{\mathrm{pinch}}=F\sin^{2}(\kappa L),$
where $L$ is the coupling length shown in Fig. 1, $\kappa$ is the coupling coefficient, and $F\leq 1$ is the maximum coupling efficiency. The factors $\kappa$ and $F$ are influenced by various factors, including the effective refractive index and the transverse modes of both the waveguide and the separate dielectric. It can be observed from the above formulas that the power emitted by a PA can be controlled by adjusting the coupling length $L$. When the waveguide and the separate dielectric share the same effective refractive index, the coupling efficiency $F$ reaches its maximum value of $1$, making it possible to radiate full power from a single PA with a coupling length of $\pi/(2\kappa)$.
@@ II-B Signal Model
As shown in Fig. 1, the composite channel from the base station (BS) to the user in the PASS system consists of two components: in-waveguide propagation and free-space propagation. Compared to free-space propagation, in-waveguide propagation incurs almost negligible path loss, e.g., approximately 0.01-0.03 dB/m for a circular copper waveguide at 15 GHz [13]. Consequently, when multiple PAs are deployed on a waveguide, the signal received at the user from the $n$-th PA through the LoS channel can be expressed mathematically as:
$y_{n}=\frac{\beta_{n}\sqrt{P_{n}}}{r_{n}}e^{-j\frac{2\pi}{\lambda}\left(r_{n}+n_{\mathrm{eff}}d_{n}\right)}x.$
In this expression, $x$ is the signal fed into the waveguide, $\beta_{n}$ is the complex channel of the free-space propagation at the reference distance of 1 meter, $P_{n}$ is the power radiated from the PA, $\lambda$ denotes the signal wavelength, $n_{\mathrm{eff}}$ represents the effective refractive index of the waveguide, $r_{n}$ is the distance of the free-space propagation from the PA to the user, and $d_{n}$ is the distance of the waveguide propagation. By adjusting the location of PAs, both $r_{n}$ and $d_{n}$ can be modified, which in turn affects the amplitude and phase of the received signal.
Based on the physical principle of PAs, the power model of PASS, signified by $P_{n}$, is significantly different from that of conventional multi-antenna systems. Specifically, with conventional antennas, the power of the signal emitted from each antenna is actively controlled by electronic devices including radio-frequency (RF) chains and power splitters. In contrast, for PAs, the transmit power is passively determined by the coupling length with the waveguide. Furthermore, when multiple PAs are pinched along a waveguide, the power radiated by subsequent antennas is influenced by the power radiated by preceding ones, leading to a coupled power model. While the coupling length of a PA is not as easily adjustable as the electronic parameters of conventional antennas since PAs are typically passive, it is still possible to design and manufacture PAs with specific coupling lengths to achieve a desired power model. In the following, we introduce two common power models for PAs.
  • Equal Power Model: In this model, the PAs on the waveguide radiates an equal amount of power. This can be achieved by designing the PAs with varying coupling lengths, such that the earlier PAs radiate a smaller proportion of the power in the waveguide, while the later PAs radiate a larger proportion of the remaining power.
@@
These two power models offer different pros and cons. Compared to the proportional power model, where the latter PAs may radiate only a smaller amount of power, the equal power model ensures that all PAs are effectively utilized. The equal power model also simplifies the performance analysis of the PASS system due to its simple mathematical formulation, which is useful to identify the fundamental limits of PASS, and hence has contributed to its widespread adoption in the literature. In contrast, the proportional power model requires less stringent manufacturing specifications, as all PAs share the same design, thus reducing production complexity and costs.
Figure 2: Non-Multiplexing transmission architectures for PASS
@@ II-C Communication System Design
From a communication system design perspective, it is crucial to position the PAs at suitable locations to maximize communication performance. On the one hand, the PAs should be deployed close to the user to mitigate the significant signal attenuation caused by free-space path loss and blockage. On the other hand, when multiple PAs are deployed along the waveguide, they form an antenna array, enabling pinching beamforming to enhance signal strength and reduce interference. By fine-tuning the positions of PAs along the waveguide, the phase shifts of the signals, due to the in-waveguide and free-space propagation, can be adjusted, thereby optimizing the beamforming gain. Leveraging the flexibility and portability of PAs, their positions can be determined either actively by the BS or passively by the user, leading to the following two designs:
  • BS-centric Design: In this design, the positions of PAs are determined by the BS to optimize specific performance metrics of the overall PASS system, such as spectral efficiency and user fairness. In dynamic scenarios, the positions of PAs can be adjusted actively based on the locations of all users in two ways: either by moving them along a pre-installed track parallel to the waveguide for continuous activation, or by selectively activating PAs at the pre-configured locations on the waveguide for discrete activation.
  • User-centric Design: In this design, users can pinch or release portable PAs onto the waveguide as needed to enhance communication service in their local area. Whenever a PA is pinched or released by a user, the baseband processing strategy needs to be optimized and updated based on the new propagation environment to avoid the performance degradation to other users.
By comparing these two designs, it can be concluded that the BS-centric design offers high flexibility, allowing the PASS to effectively mitigate path loss and blockage issues while enhancing beamforming gain for all users. However, this comes with a higher implementation cost. In contrast, the user-centric design has a lower implementation cost but may result in each PA being effectively utilized for only a single user.
@@ III Transmission Architectures for PASS
III-A Non-Multiplexing Architecture
A schematic illustration of the non-multiplexing architecture is shown in Fig. 2. In this architecture, the data streams to be transmitted are not multiplexed at the baseband, meaning that each waveguide is fed individually by a single data stream. In this setup, the primary role of the baseband processing is to manage the power allocation across these data streams to optimize transmission efficiency. The key benefit of this design is to reduces implementation complexity. Compared to the conventional multi-antenna systems, the unique capability of pinching beamforming, achieved by adjusting the positions of PAs, can be utilized to enhance communication performance. This non-multiplexing architecture is particularly well-suited for those commonly encountered scenarios, where each waveguide serves geographically isolated areas, minimizing the need for advanced interference management between waveguides.
Figure 3: Multiplexing transmission architectures for PASS
@@ III-A Non-Multiplexing Architecture
III-B Multiplexing Architecture
As shown at the bottom of Fig. 3, the multiplexing architecture multiplexes multiple data streams before feeding them into the waveguides, thus facilitating the joint transmit and pinching beamforming. This architecture can be seen as an extension of the conventional MIMO transmission architecture, where each electronic antenna is replaced by a waveguide. The key difference between the two architectures lies in the point of signal radiation. In conventional MIMO systems, signals are radiated directly from the electronic antennas into free space. In contrast, in the multiplexing architecture of PASS, signals are radiated at specific locations along the waveguide via PAs. Consequently, unlike conventional MIMO, which relies on a massive number of electronic antennas to achieve high spatial degrees of freedom (DoFs) [14], the multiplexing architecture of PASS significantly reduces the requirement for numerous waveguides because a large number of low-cost PAs can be easily installed on a single waveguide. Based on the above discussion, the multiplexing architecture for PASS can be designed in analogy to conventional MIMO architecture but with distinct differences in implementation and design focus. In the following, we introduce three basic multiplexing architectures, as shown in Fig. 3, namely sub-connected architecture, fully-connected architecture, and phase shifter (PS)-based fully-connected architecture.

### v3

@@ Abstract
Flexible-antenna systems have recently attracted significant research attention due to their potential to intelligently reconfigure wireless channels. However, the current flexible-antenna systems still suffer from fundamental limitations, such as free-space path loss and line-of-sight blockage. This article introduces a novel flexible-antenna system, termed the Pinching-Antenna SyStem (PASS). PASS adopts the dielectric waveguides as the primary transmission medium and radiates signals into free space by flexibly pinching discrete dielectric particles, referred to as pinching antennas, along the waveguide. By combining the strengths of both wireless and wired communication, PASS effectively mitigates inherent wireless limitations while offering high antenna reconfigurability. This article reviews the key features of PASS in comparison with conventional wireless systems, analyzes its main advantages, and discusses potential designs, transmission architectures, and application scenarios. Finally, it outlines promising research directions and open challenges associated with PASS.
@@ I Introduction
Building on this momentum, recent advancements in intelligent and flexible antenna designs have further introduced novel approaches for channel reconfiguration. Technologies, such as reconfigurable intelligent surfaces [3, 4], dynamic metasurface antennas [5], fluid antennas [6], and movable antennas [7], can dynamically adjust their positions, apertures, and electromagnetic properties, enabling precise control of wireless channel conditions. By optimizing their reconfigurable parameters, these novel antenna types and array architectures offer flexibility in controlling the effective end-to-end channel gain. This capability has demonstrated great potential for improving system performance and has attracted considerable research interest.
Despite their potential to improve network throughput, current flexible-antenna systems, in which antenna reconfigurations are often confined to small apertures, face two fundamental wireless limitations: free-space path loss and line-of-sight (Los) blockage. Free-space path loss leads to rapid signal attenuation, particularly over long distances or at high frequencies, while LoS blockage can cause deep fading and create coverage dead zones. To overcome these challenges, a novel, highly flexible, and efficient design, known as the Pinching-Antenna SyStem (PASS) [8, 9], has been recently proposed.
@@
I-A The Road to Near-Wired Connectivity: A Prototype of PASS
The most effective way to overcome free-space path loss and LoS blockage is through cable-based wired communication, which completely eliminates both issues. However, wired links are inherently unsuitable for mobile scenarios. This limitation raises a fundamental question: Can we bridge the gap between fixed cables and fully over-the-air communication? A promising answer is PASS, first developed and prototyped by NTT DOCOMO in 2021 [8], as illustrated in Fig. 1. PASS is based on low-loss dielectric waveguides that serve both as a transmission medium and as an “antenna rail.” At an arbitrary point along the waveguide, a small dielectric particle can be pinched onto the waveguide, like attaching a clothespin to a line, to form a pinching antenna (PA). Each PA functions as a controlled leakage point, enabling the guided wave to radiate into or receive energy from free space. This makes PASS act as a reconfigurable waveguide and provides greater flexibility than traditional leaky-wave antennas.
Unlike conventional flexible-antenna systems, a PASS waveguide can be run for tens of meters or even longer, allowing PAs to be positioned within LoS of users [8, 9]. This adaptability effectively creates a near-wired connection, ensuring favorable channel conditions and shifting the focus of wireless communication networks from the “last mile” to the more manageable “last meter”. Therefore, PASS combines the high reliability of wired communication with the deployment flexibility of wireless access, offering a novel and effective approach to overcoming the limitations of conventional wireless systems. This unique capability is reflected in the name “PASS”, highlighting its ability to provide the shortest and most efficient wireless connections between the PAs and the wireless transceivers, which effectively pass wireless signals with minimal loss.
@@ I-B Key Advantages for Employing PASS in Wireless Communications
  • Enhanced LoS Stability and “Last Meter” Communications: By extending the waveguide and enabling the flexible deployment of PAs, PASS effectively mitigate large-scale path loss and establishes stable LoS communication environments. This capability enhances “last meter” communications and expands network coverage, which is not achievable with other flexible-antenna technologies.
@@
  • Cost-Effective and Scalable Implementation: PASS are simple to implement and economical to deploy, not only because of the low cost of waveguides, but also because PASS only require adding or removing dielectric materials at selected points along the waveguide [8]. Furthermore, PASS theoretically allows the addition or removal of any number of radiating units along the waveguide in a low-cost manner, making them a scalable and flexible solution for practical network design.
These advantages have fueled growing research interest in PASS. The performance of PASS has been analyzed in terms of its ergodic rate [9] and its array gain [11]. Furthermore, various activation algorithms have been proposed to optimize the locations of PAs along a single waveguide, such as [9] and [12]. However, despite these initial efforts, the design of PASS for wireless communications remains in its early stage. Therefore, this article aims to provide a concise yet comprehensive overview of existing PASS techniques and to offer new insights into their development, serving as a useful reference for researchers interested in exploring the vast potential of PASS. We begin by discussing the fundamental differences between PASS and conventional multi-antenna systems. Next, we introduce several potential transmission architectures to support the design of pinching beamforming. We then explore promising application scenarios and evaluate the corresponding system performance. Finally, we outline key open challenges to stimulate further research in this emerging area.
@@ II Pinching Antenna Fundamentals:
In this section, we introduce the fundamentals of PAs and highlight the basic differences from conventional electronic antennas from their physical principle, signal model, and communication system design perspective, respectively.
@@ II-A Physical Principle
Unlike conventional electronic antennas, which rely on electronic devices to convert alternating electric currents into electromagnetic (EM) waves, the physical principle of PAs is based on the phenomenon of EM coupling that arises when a separate dielectric is brought into close proximity with a dielectric waveguide [8]. This coupling enables the power exchange between the waveguide and the separate dielectric, as illustrated in the bottom-right of Fig. 1. Therefore, by pinching a separate dielectric onto the waveguide, the EM waves originally confined within the waveguide are coupled into the separate dielectric, where they are subsequently radiated into free space, thereby forming an “antenna”. Unlike conventional electronic antennas, PAs are entirely passive, simpler to manufacture and deploy, and hence provide greater position flexibility.
@@
$P_{\mathrm{guide}}$ $=1-F\sin^{2}(\kappa L),\quad P_{\mathrm{pinch}}=F\sin^{2}(\kappa L),$ (1)
where $L$ is the coupling length shown in Fig. 1, $\kappa$ is the coupling coefficient, and $F\leq 1$ is the maximum coupling efficiency. When the waveguide and the separate dielectric share the same effective refractive index, the coupling efficiency $F$ reaches its maximum value of $1$, making it possible to radiate full power from a single PA with a coupling length of $\pi/(2\kappa)$. Moreover, from the above formula, it can be observed that the power emitted by each PA can be controlled by adjusting the coupling length $L$ and the coupling coefficient $\kappa$. The coefficients $\kappa$ and $F$ are influenced by various factors, including the effective refractive index, transverse modes, and the spacing between the waveguide and the separate dielectric.
@@ II-B Signal Model
As shown in Fig. 1, the composite channel from the base station (BS) to the user in the PASS system consists of two components: in-waveguide propagation and free-space propagation. Compared to free-space propagation, in-waveguide propagation incurs almost negligible path loss. For example, a ceramic-ribbon dielectric waveguide operating in the millimeter-submillimeter band attenuates by less than 0.01 dB/m [14]. Consequently, when multiple PAs are deployed on a waveguide, the signal received at the user from the $n$-th PA through the LoS channel can be modeled as:
$y_{n}=\frac{\beta_{n}\sqrt{P_{n}}}{r_{n}}e^{-j\frac{2\pi}{\lambda}\left(r_{n}+n_{\mathrm{eff}}d_{n}\right)}x.$ (2)
In this expression, $j=\sqrt{-1}$, $x$ is the signal fed into the waveguide, $\beta_{n}$ is the complex channel of the free-space propagation at the reference distance of 1 meter, $P_{n}$ is the power radiated from the PA, $\lambda$ denotes the signal wavelength, $n_{\mathrm{eff}}$ represents the effective refractive index of the waveguide, $r_{n}$ is the distance of the free-space propagation from the PA to the user, and $d_{n}$ is the distance of the waveguide propagation. By adjusting the location of PAs, both $r_{n}$ and $d_{n}$ can be modified, which in turn affects the amplitude and phase of the received signal.
Based on the physical principle of PAs, the power model of PASS, signified by $P_{n}$, is significantly different from that of conventional multi-antenna systems. Specifically, with conventional antennas, the power of the signal emitted from each antenna is actively controlled by electronic devices, including radio-frequency (RF) chains and power splitters. In contrast, for PAs, the transmit power is passively determined by the coupling length with the waveguide. Furthermore, when multiple PAs are pinched along a waveguide, the power radiated by subsequent antennas is influenced by the power radiated by preceding ones, leading to a coupled power model. While the coupling length of a PA is not as easily adjustable as the electronic parameters of conventional antennas, since PAs are typically passive, it is still possible to design and manufacture PAs with specific coupling lengths to achieve a desired power model. In the following, we introduce two common power models for PAs.
  • Equal Power Model: In this model, the PAs on the waveguide radiate an equal amount of power. This can be achieved by designing the PAs with varying coupling lengths, such that the earlier PAs radiate a smaller proportion of the power in the waveguide, while the later PAs radiate a larger proportion of the remaining power.
@@
These two power models offer different pros and cons. Compared to the proportional power model, where the latter PAs may radiate only a smaller amount of power, the equal power model ensures that all PAs are effectively utilized. The equal power model also simplifies the performance analysis of the PASS system due to its simple mathematical formulation, which is useful to identify the fundamental limits of PASS, and therefore has contributed to its widespread adoption in the literature. In contrast, the proportional power model requires less stringent manufacturing specifications, as all PAs share the same design, thus reducing production complexity and costs.
Figure 2: Waveguide division architectures for PASS
@@ II-C Communication System Design
From a communication system design perspective, it is crucial to position the PAs at suitable locations to maximize communication performance. On the one hand, the PAs should be deployed close to the user to mitigate the significant signal attenuation caused by free-space path loss and blockage. On the other hand, when multiple PAs are deployed along the waveguide, they form an antenna array, enabling pinching beamforming to enhance signal strength and reduce interference. Note that pinching beamforming poses greater challenges compared to existing flexible-antenna systems. This is because fine-tuning the positions of the PAs along the waveguide enables simultaneous adjustment of the phase shifts of in-waveguide and free-space propagation, as well as the free-space path loss, to achieve the desired beamforming gain. Leveraging the flexibility and portability of PAs, their positions can be determined either actively by the BS or passively by the user, leading to the following two designs:
  • BS-centric Design: In this design, the positions of PAs are determined by the BS to optimize specific performance metrics of the overall PASS system, such as spectral efficiency and user fairness. In dynamic scenarios, there are two typical PA activation schemes, namely continuous activation and discrete activation. For continuous activation, a slide track is installed parallel to the waveguide, allowing the attached PAs to be mechanically repositioned along the waveguide, e.g., using micro motors. For discrete activation, a large number of PAs are pre-installed at various locations along the waveguide and dynamically activated through fast-response electrical means, such as using electromagnetic relays. It is worth noting that mechanical continuous activation may introduce considerable time delays, making it less efficient for fast-varying wireless channels compared to electrical discrete activation. However, it offers more precise PA positioning and is preferable for slow-varying wireless environments.
  • User-centric Design: In this design, users can pinch or release portable PAs onto the waveguide as needed to enhance communication service in their local area. Whenever a PA is pinched or released by a user, the baseband processing strategy needs to be optimized and updated based on the new propagation environment to avoid performance degradation for other users.
By comparing these two designs, it can be concluded that the BS-centric design offers high flexibility, allowing the PASS to effectively mitigate path loss and blockage issues while enhancing beamforming gain for all users. However, this comes with a higher implementation cost. In contrast, the user-centric design has a lower implementation cost but may result in each PA being effectively utilized for a single user.
@@ III Transmission Architectures for PASS
III-A Waveguide Division Architecture
A schematic illustration of the waveguide division architecture is shown in Fig. 2. In this architecture, the data streams to be transmitted are not multiplexed at the baseband, meaning that each waveguide is fed individually by a single data stream. In this setup, the primary role of the baseband processing is to manage the power allocation across these data streams to optimize transmission efficiency. The key benefit of this design is to reduce implementation complexity. Compared to the conventional multi-antenna systems, the unique capability of pinching beamforming, achieved by adjusting the positions of PAs, can be utilized to enhance communication performance. This waveguide division architecture is particularly well-suited for those commonly encountered scenarios, where each waveguide serves geographically isolated areas, minimizing the need for advanced interference management between waveguides.
Figure 3: Waveguide multiplexing architectures for PASS
@@ III-A Waveguide Division Architecture
III-B Waveguide Multiplexing Architecture
As shown at the bottom of Fig. 3, the waveguide multiplexing architecture multiplexes multiple data streams before feeding them into the waveguides, thus facilitating the joint transmit and pinching beamforming. This architecture can be seen as an extension of the conventional MIMO transmission architecture, where each electronic antenna is replaced by a waveguide. The key difference between the two architectures lies in the point of signal radiation. In conventional MIMO systems, signals are radiated directly from the electronic antennas into free space. In contrast, in the waveguide multiplexing architecture of PASS, signals are radiated at specific locations along the waveguide via PAs. Consequently, unlike conventional MIMO, which relies on a massive number of electronic antennas to achieve high spatial degrees of freedom (DoFs) [15], the waveguide multiplexing architecture significantly reduces the requirement for numerous waveguides because a large number of low-cost PAs can be easily installed on a single waveguide. Based on the above discussion, the waveguide multiplexing architecture can be designed in analogy to conventional MIMO architecture, but with distinct differences in implementation and design focus. In the following, we introduce three basic waveguide multiplexing architectures, as shown in Fig. 3, namely sub-connected architecture, fully-connected architecture, and phase shifter (PS)-based fully-connected architecture.